Grade 10Mechanics

the dimensional formula of a physical quantity is given as [M-1L3T-2].the errors in measuring M,L and T respectively are 2%,3% and 4% .the maximum percentage of error that occurs in measuring the quantity X is?

meenu s nair
8 Years agoGrade 10

1 Answer

Arun
8 Years ago
Dear student
 
Maximum percentage of error in that physical quantity =  1 * 2  + 3 * 3 + 2 * 4 = 2 + 9 + 8 = 19%
hope it helps
 
Regards
Arun (askIITians forum expert)